OpenStudy (anonymous):

When solving this equation by completing the square, which of the following shows the equation with the square completed? x2 - 4x - 32 = 0 A. (x - 4)2 = 36 B. (x - 2)2 = 36 C. (x - 8)2 = 4 When solving this equation by completing the square, which of the following shows the equation with the square completed? x2 - 4x - 32 = 0 A. (x - 4)2 = 36 B. (x - 2)2 = 36 C. (x - 8)2 = 4 @Mathematics

OpenStudy (anonymous):

I believe there's an x missing somewhere.

OpenStudy (anonymous):

\[x^2 - 4x - 32 = 0\] \[x^2 -4x + 4 - 32 = 4\] \[(x-2)^2 = 36\] therefore the right answer is B

OpenStudy (anonymous):

Ohhh that's a squared, not a 2 lol